OpenStudy (anonymous):

Help would be much appreciated: A 240.3 kg object is located between the earth and the moon. The mass of the earth is 6.0x10^24 kg and the mass of the moon is 7.40x10^22 kg. The distance from the earth's center to the moon's center is 3.84x10^8meters. The distance from the object to the moon's center is 2.2 times longer than the distance from the object to the earth's center. What is the net force on the object between the earth and moon?

OpenStudy (loser66):

 @douglaswinslowcooper since you are here, please give him your hand

OpenStudy (loser66):

and I need your help here, too. Please, http://openstudy.com/study#/updates/529ea5c3e4b030bcac5de621

OpenStudy (anonymous):

Start with F = -GmM/r^2, where m is the mass of the object, M is the mass of the earth, and we will also use M' as the mass of the moon. We can get the distances to use for r from the information that gives us R from the earth to the moon, and that R = R1 + R2, where R2 = 2.2 R1, so R1 + 2.2 R1 = 3.2 R1 = R and R1 = R/3.2, R2 = (2.2/3.2) R The force on the object will be the difference between the force on it from the moon and the force on it from the earth: net force = moon - earth = m M' G/R2^2 - m M G/R1^2 The rest, as is said, is "an exercise for the student." @Loser96 Thanks for complimentary invitation and helpful diagram.

OpenStudy (anonymous):

thank you so much!!!!

OpenStudy (anonymous):

you guys rock! makes a lot more sense now.